A boy is swinging on playground exercise gym. He swings through an arc of 1 m on his first swing. On each successive swing, he swings through an arc 70% of the previous length.

a) Show that the total distance the boy swings in his first n swings is 10

given by S  3 1 0.7 where n 1,2,3,.... nn

Hence, find the total distance travelled by the boy after six swings.


1
Expert's answer
2021-11-10T17:27:47-0500

Let "a_n"  denote the distance the boy swings in the  "n^{th}" swing.

Then, according to the given information, the sequence "a_n"  is a geometric sequence with first term  "a_n=1" and common ratio "r=70\%=0.7"

Hence, the general term is given by "a_n.r^{n-1}=1.07^{n-1}=0.7^{n-1}"

Consequently, the total distance the boy swings in his first n swings is as follows:

"S_n=a_1+a_2+a_3.......+a_n\\=1+0.7^1+0.7^2+0.7^3+........+0.7^{n-1}\\=0.7s_n=0.7+0.7^2+0.7^3+......+0.7^n\\s_n-0.7 s_n=(1+0.7^1+0.7^2+0.7^3+......+0.7^{n-1}-(0.7+0.7^2+0.7^3+....+0.7^n)\\=0.3s_n=1-0.7^n\\=\frac{3}{10}s_n=1-0.7^n\\s_n=\frac{10}{3}(1-0.7^n)"

Therefore, the total distance the boy swings in his first  n swings is given by "S_n=\frac{10}{3}(1-0.7^n)"  where "n=1,2,3"

Further,  "S_6=\frac{10}{3}(1-0.7^6)=2.94"

Therefore, the total distance travelled by the boy after six swings is approximately 2.94 m.